# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ 2.1 Microscopic Boltzmann weight
Here, $N_{i}^{\#}-1$ takes care of the self-interaction of the particle. $\beta$ is the inverse temperature. We fix the Boltzmann constant to unity throughout this paper. $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ correspond to the actio = reactio symmetry preservation and violation terms, respectively. They quantify the total reciprocal and nonreciprocal interaction energy experienced by the type $i$ particle owing to all other particles. $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ are locally defined properties of each particle.
We examine the possibility of representing the locally defined microscopic Boltzmann weight as a global property of the lattice, namely the total interaction energy $\varepsilon^{int}$. The condition for it turns out to be $v_{ij}=v_{ji}$, which is satisfied by $v_{ij}^{r}$ and violated by $v_{ij}^{nr}$. Hence, the reciprocal part of the microscopic Boltzmann weight is written as the microscopic interaction energy functional $\varepsilon^{int}$ of the lattice,
@@
In the absence of non-reciprocity, i.e. $v_{ij}^{nr}=0,\forall i,j$, the system satisfies the Boltzmann distribution generated by $\varepsilon^{int}$. A detailed discussion on thermodynamics is subsequently included.
@@
The eq. 2 is not necessarily the equilibrium interaction energy functional. Consider the decomposition, $v_{ij}=v_{ij}^{eq}+v_{ij}^{neq}$. Where $v_{ij}^{eq}$ is derived from the equilibrium energy functional of the form eq. 2 and $v_{ij}^{neq}$ is the remaining microscopic non-equilibrium interaction energy attributed to the violation of Newton’s third law of motion. This leads to $v_{ij}^{r}=v_{ij}^{eq}+(v_{ij}^{neq}+v_{ji}^{neq})/2$ and $v_{ij}^{nr}=(v_{ij}^{neq}-v_{ji}^{neq})/2$, hence, $v_{ij}^{r}$ and $v_{ij}^{nr}$ incorporates the symmetric and anti-symmetric parts respectively of the $v_{ij}^{neq}$. In conclusion, even though the microscopic physical interaction rules are governed by $v_{ij},v_{ji},v_{ij}^{eq},v_{ij}^{neq}$, the underlying geometric and thermodynamic structure is governed by $v_{ij}^{r},v_{ij}^{nr}$, and $|\varepsilon^{int}|\geq|\varepsilon^{eq}|$ gives a tighter bound on the conservative interaction energy functional.
@@
Moreover, this analysis reveals that the decomposition of $v_{ij}$ and $v_{ji}$ (gauge fixing) is not necessarily unique, unless ‘the orthogonal reference gauge’ is imposed. Further, it will be shown that only ‘the orthogonal reference gauge’ leads to observation scale invariant decomposition of actio-reactio symmetry conserving and breaking terms. Consequently, it also gives the physically correct thermodynamic dissipation of the system. The decomposition of $\epsilon_{i}^{\#}$ into $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ is equivalent to the Helmholtz-Hodge decomposition of $\epsilon_{i}^{\#}$ into a conservative gradient force $\epsilon_{i}^{r}$ derived from $\varepsilon^{int}$ and a divergence-free non-conservative force $f_{i}^{nr}$.
@@ 2.2 Dynamics
Figure 2: (a) The illustration for the microscopic interaction between the particle types denoted by $v_{ij}$ confined to the lattice site ##\## such that $v_{ij}\neq v_{ji}$. (b) All possible transitions are denoted by a curved harpoon. The exact expression for two transition rates $k_{\gamma\gamma^{\prime}}$ in terms of $v_{ij}$ for the microscopic lattice configuration in fig. 1(a) are demonstrated.
@@ A Single particle microstate transition rates
The particle can change its type or the lattice location, which is defined as the reactive and diffusive transition respectively. A reactive transition from type $\gamma^{\prime}$ to $\gamma$ in ##\## is indicated by $\Delta_{\gamma\gamma^{\prime}}^{\#}$ resulting in $\{N\}\to\{N+\Delta_{\gamma\gamma^{\prime}}^{\#}\}$. Similarly, a diffusive transition of type $i$ particle from ##\## in direction $\vec{\mathcal{D}}$ is indicated by $\Delta_{i}^{\vec{\mathcal{D}}\#}$ resulting in $\{N\}\to\{N+\Delta_{i}^{\vec{\mathcal{D}}\#}\}$. The transition rate of a reactive ($k_{\gamma^{\prime}\gamma}^{\#}$) and diffusive ($k_{i}^{\vec{\mathcal{D}}\#}$) jump are constructed using $\epsilon_{i}^{\#}$:
@@
The LDB for the forward and backward microscopic transitions read:
@@
The environment supports the thermodynamic cost for the change in the microscopic Boltzmann weight. We have modified eq. 4 by introducing a non-conservative external driving force $f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{f}_{i}^{sp}$ that correspond to chemical and self-propulsion driving. The thermodynamic cost for such an external driving is supported by an externally coupled reservoir of the enzymatic particles (for the reactive transition) or colloidal particles (for the diffusive transitions) respectively. We use shorthand notations $\Delta_{\gamma\gamma^{\prime}}^{\#}(*)=(*)_{\gamma^{\prime}}^{\#}-(*)_{\gamma}^{\#}$ and $\Delta_{i}^{\vec{\mathcal{D}}\#}(*)=(*)_{i}^{\#}-(*)_{i}^{\vec{\mathcal{D}}\#}$ for the change in $\epsilon_{i}^{\#},\epsilon_{i}^{r}$ and $f_{i}^{nr}$ due to a transition. We introduce a shorthand notation $\Delta$ for referring to a transition, applied as a diffusive or reactive transition depending on the context. The set of all reactive transitions is denoted by $\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}$. The set of all diffusive transitions is denoted by $\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}$. Four such possibilities for a two-dimensional square lattice exist: upward, downward, leftward, and rightward. The microscopic interactions and transitions are shown in fig. 2.
B Multi-particle microstate transition rates
@@
Particles with identical type and lattice index have the same dynamic and thermodynamic properties. Hence, $k_{\gamma\gamma^{\prime}}^{\#}$ are lumped together to give multi-particle transition rates $k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})=N_{\gamma^{\prime}}^{\#}k_{\gamma\gamma^{\prime}}^{\#}$ and ${k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})}=N_{i}^{\#}k_{i}^{\vec{\mathcal{D}}\#}$. The LDB for the multi-particle transition rate for the reactive (${k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})}$) and diffusive (${k_{{i}}^{\vec{\mathcal{D}}\#}(\{N\})}$) is:
@@
$\begin{split}\frac{k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})}{k_{\gamma^{\prime}\gamma}^{\#}({\{N+\Delta_{\gamma\gamma^{\prime}}^{\#}\}})}&=e^{-\Delta_{\gamma\gamma^{\prime}}^{\#}s^{b}+\Delta_{\gamma\gamma^{\prime}}^{\#}\epsilon+f_{\gamma\gamma^{\prime}}^{ch}}.\\ \frac{k_{{i}}^{\vec{\mathcal{D}}\#}(\{N\})}{k_{i}^{(\vec{\mathcal{D}}\#)^{-1}}(\{N+\Delta_{i}^{\vec{\mathcal{D}}\#}\})}&=e^{-\Delta_{i}^{\vec{\mathcal{D}}\#}s^{b}+\Delta_{i}^{\vec{\mathcal{D}}\#}\epsilon+\vec{\mathcal{D}}\cdot\vec{f}_{i}^{sp}}.\end{split}$ (5)
@@
Where, $\Delta s^{b}$ incorporates the change in microscopic Boltzmann entropy defined as a global lattice property [91]:
@@
It quantifies the statistical degeneracy of the microstate and is a system property defined for the lattice. Importantly, $\Delta_{\gamma\gamma^{\prime}}^{\#}\epsilon^{r}=\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}$ and $\Delta_{i}^{\vec{\mathcal{D}}\#}\epsilon^{r}=\Delta_{i}^{\vec{\mathcal{D}}\#}\varepsilon^{int}$ is satisfied by the reciprocal microscopic Boltzmann weight. We introduce notation $\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}=\varepsilon^{int}(\{N\})-\varepsilon^{int}(\{N+\Delta_{\gamma\gamma^{\prime}}\})$, which is resolved as a change of a function of the lattice state, a global property, this applies similarly to $s^{b}$. In contrast, the change in non-reciprocal microscopic Boltzmann weight $\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$ can not be represented as a change of a global function of the lattice state. This results in a fundamental difference between the boundary ($\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}$ and $-\Delta_{\gamma\gamma^{\prime}}^{\#}s^{b}$) and the bulk terms ($\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$) of the transition affinity.

### v3

@@ A Theoretical formulation
We drop the lattice index #\# for $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ for brevity, but it is understood that these quantities are evaluated with the appropriate lattice #\# and type $i$ indexes. Here, $N_{i}^{\#}-1$ accounts for the self-interaction of the particle. $\beta$ is the inverse temperature, and the Boltzmann constant is set to unity throughout this work. $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ correspond to the actio=reactio symmetry preserving and breaking terms, respectively, and quantify the total reciprocal and non-reciprocal interaction energy experienced by the type $i$ particle due to all other particles. $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ are local properties of each particle. For non-interacting ideal particles, $\epsilon_{i}^{\#}=0$, since $v_{ij}=0,\forall j$. Table 1 summarizes the notation for microscopic systems.
We examine the possibility of representing the locally defined microscopic Boltzmann weight as a global property of the lattice, namely the total interaction energy $\varepsilon^{int}$. This is possible only if $v_{ij}=v_{ji}$, which is satisfied by $v_{ij}^{r}$ and violated by $v_{ij}^{nr}$. Hence, the reciprocal part of the microscopic Boltzmann weight can be expressed as the microscopic interaction energy functional $\varepsilon^{int}$ of the lattice,
@@
In the absence of non-reciprocity, i.e., $v_{ij}^{nr}=0,\forall i,j$, the system satisfies the Boltzmann distribution generated by $\varepsilon^{int}$. A detailed discussion on thermodynamics follows.
@@
Equation eq. 2 does not necessarily represent the equilibrium interaction energy functional. Consider a decomposition $v_{ij}=v_{ij}^{eq}+v_{ij}^{neq}$, where $v_{ij}^{eq}$ is derived from the equilibrium energy functional of the form eq. 2 and $v_{ij}^{neq}$ is the remaining microscopic non-equilibrium interaction energy attributed to the violation of Newton’s third law. This gives $v_{ij}^{r}=v_{ij}^{eq}+(v_{ij}^{neq}+v_{ji}^{neq})/2$ and $v_{ij}^{nr}=(v_{ij}^{neq}-v_{ji}^{neq})/2$, so that $v_{ij}^{r}$ and $v_{ij}^{nr}$ incorporate the symmetric and anti-symmetric parts of $v_{ij}^{neq}$, respectively. In conclusion, even though the microscopic physical interaction rules are governed by $v_{ij},v_{ji},v_{ij}^{eq},v_{ij}^{neq}$, the underlying geometric and thermodynamic structure is captured by $v_{ij}^{r},v_{ij}^{nr}$, and $|\varepsilon^{int}|\geq|\varepsilon^{eq}|$ gives a tighter bound on the conservative interaction energy functional.
@@
Moreover, this analysis shows that the decomposition of $v_{ij}$ and $v_{ji}$ (gauge fixing) is not necessarily unique, unless ‘the orthogonal reference gauge’ is imposed. Further, only ‘the orthogonal reference gauge’ leads to an observation scale-invariant decomposition of actio-reactio symmetry conserving and breaking terms. Consequently, it also yields the physically correct thermodynamic dissipation of the system. The decomposition of $\epsilon_{i}^{\#}$ into $\epsilon_{i}^{r}$ and $f_{i}^{nr}$ is equivalent to the Helmholtz-Hodge decomposition of $\epsilon_{i}^{\#}$ into a conservative gradient force $\epsilon_{i}^{r}$ derived from $\varepsilon^{int}$ and a divergence-free non-conservative force $f_{i}^{nr}$.
Table 1: Summary of notation used for microscopic interacting particles. Throughout this work, superscript #\# and subscript $i$ denote the lattice site and particle-type index, respectively. For microscopic interaction coefficients between particle types, two indices are needed; the convention ‘left–right’ denotes ‘influenced–influencer’ particle. Reactive or diffusive transitions correspond to changes in particle type or lattice location, respectively. For transitions, the ‘right-to-left’ notation denotes ‘initial-to-final’ microstate. Latin letters $i,j$ refer to state-space properties requiring the particle type specification, e.g., $v_{ij},v_{ij}^{r},v_{ij}^{nr},\epsilon_{i}^{\#},\epsilon_{i}^{r},f_{i}^{nr}$. Greek letters $\gamma,\gamma^{\prime}$ refer to transition-space properties specifying particle type changes, e.g., $\Delta_{\gamma\gamma{{}^{\prime}}}^{\#},f_{\gamma\gamma{{}^{\prime}}}^{ch},k_{\gamma\gamma{{}^{\prime}}}^{\#}$. This distinction avoids confusion between state-space and transition-space quantities.
Notation Significance
$v_{ij}$ Microscopic interaction potential energy experienced by type $i$ particle due to the presence of a type $j$ particle.
$v_{ij}^{r}$ Reciprocal component of $v_{ij}$ satisfying ‘actio=reactio’ symmetry: $v_{ij}^{r}=v_{ji}^{r}$.
$v_{ij}^{nr}$ Non-reciprocal component of $v_{ij}$ satisfying ‘actio=-reactio’ anti-symmetry: $v_{ij}^{nr}=-v_{ji}^{nr}$.
$\epsilon_{i}^{\#}$ Microscopic Boltzmann weight of a type $i$ particle at lattice site #\#, quantifying the total thermodynamic cost of placing the particle at the site.
$\epsilon_{i}^{r}$ Reciprocal part of $\epsilon_{i}^{\#}$ that respects ‘actio=reactio’ symmetry.
$f^{nr}_{i}$ Non-reciprocal part of $\epsilon_{i}^{\#}$ that violates ‘actio=reactio’ symmetry.
$\varepsilon^{int}$ Total reciprocal interaction energy for the lattice.
$\Delta_{\gamma\gamma^{\prime}}^{\#}$ Reactive transition of type $\gamma^{\prime}$ particle at lattice site #\# to type $\gamma$, changing lattice configuration $\{N\}\to\{N+\Delta_{\gamma\gamma^{\prime}}^{\#}\}$.
$f_{\gamma\gamma^{\prime}}^{ch}$ External driving force along the transition $\Delta_{\gamma\gamma^{\prime}}^{\#}$, supported by an enzymatic chemical reservoir.
$k_{\gamma\gamma^{\prime}}^{\#}$ Reactive transition rate corresponding to $\Delta_{\gamma\gamma^{\prime}}^{\#}$.
$\Delta_{i}^{\vec{\mathcal{D}}\#}$ Diffusive transition of type $i$ particle at lattice site #\# in direction $\vec{\mathcal{D}}$, changing lattice configuration $\{N\}\to\{N+\Delta_{i}^{\vec{\mathcal{D}}\#}\}$.
$\vec{f}_{i}^{sp}$ Self-propulsion driving force along the transition $\Delta_{i}^{\vec{\mathcal{D}}\#}$, supported by a colloidal reservoir.
$k_{i}^{\vec{\mathcal{D}}\#}$ Diffusive transition rate corresponding to $\Delta_{i}^{\vec{\mathcal{D}}\#}$.
B Illustration of the model setup
Figure 2: (a) Illustration of the microscopic interactions between particle types, denoted by $v_{ij}$, confined to the lattice site #\#, with $v_{ij}\neq v_{ji}$. (b) All possible transitions are indicated by curved harpoons. The exact expressions for two transition rates $k_{\gamma\gamma^{\prime}}$ in terms of $v_{ij}$ for the microscopic lattice configuration in fig. 1(a) are demonstrated.
We illustrate the significance of the concept of Boltzman weights in non-reciprocal systems with a minimal model, namely the sheep–dog dynamics. Here, $v_{sd}$ represents the energy cost for an external agent to place a sheep on the lattice when a dog is already present. Since the sheep repels the dog, let us assume that 5 units of energy are required to implement this operation experimentally, implying $v_{sd}=5$ units. Similarly, $v_{ds}$ is the energy cost for an external agent to place a dog on the lattice when a sheep is already present. Because the dog is attracted to the sheep, this operation releases energy. This calibration is equivalently achieved by performing the inverse operation of removing the dog from the lattice when both sheep and dog are present, which requires $v_{ds}$ units of energy (note the sign convention due to defining the cost for the inverse operation). Let us assume that 3 units of energy are needed to remove the dog from the lattice in the presence of the sheep. Therefore, $v_{ds}=-3$ units, signifying an attractive interaction of the dog towards the sheep. All parameters of our theoretical model, namely the microscopic interaction energies $v_{ij}$ for different particle types, are to be interpreted in this way.
@@ A Single particle microstate transition rates
The particle can change its type or its lattice location, which are defined as the reactive and diffusive transitions, respectively. A reactive transition from type $\gamma^{\prime}$ to $\gamma$ at the lattice site #\# is indicated by $\Delta_{\gamma\gamma^{\prime}}^{\#}$, resulting in $\{N\}\to\{N+\Delta_{\gamma\gamma^{\prime}}^{\#}\}$, see fig. 2(b). Similarly, a diffusive transition of a type $i$ particle from the lattice site #\# in the direction $\vec{\mathcal{D}}$ is indicated by $\Delta_{i}^{\vec{\mathcal{D}}\#}$, resulting in $\{N\}\to\{N+\Delta_{i}^{\vec{\mathcal{D}}\#}\}$. The transition rates of reactive ($k_{\gamma^{\prime}\gamma}^{\#}$) and diffusive ($k_{i}^{\vec{\mathcal{D}}\#}$) jumps are constructed using $\epsilon_{i}^{\#}$:
@@
The thermodynamically consistent exponential parameterization of $k_{\gamma^{\prime}\gamma}^{\#}$ and $k_{i}^{\vec{\mathcal{D}}\#}$ in $\epsilon_{i}^{\#}$ allows one to visualize the thermodynamic meaning of $\epsilon_{i}^{\#}$. For non-interacting ideal particles, $\epsilon_{i}^{\#}=0$; therefore, $k_{\gamma^{\prime}\gamma}^{\#}=d_{\gamma^{\prime}\gamma}$ and $k_{i}^{\vec{\mathcal{D}}\#}=d_{i}^{\mathcal{D}}$ are the equilibrium transition rates for non-interacting ideal particles, which also quantify the strength of the transition rates in eq. 3. They satisfy $d_{\gamma\gamma^{\prime}}=d_{\gamma^{\prime}\gamma}$ and $d_{i}^{\mathcal{D}}=d_{i}^{\mathcal{D}^{-1}}$.
The LDB condition for the forward and backward microscopic transitions constructed using eq. 3 reads:
@@
Equation 4 expresses the microscopic LDB for non-reciprocal particles in a more familiar form compared to its original definition in eq. 3. However, the time-symmetric part of the transition rates cancels out in eq. 4 due to the relations $d_{\gamma^{\prime}\gamma}=d_{\gamma\gamma^{\prime}}$ and $d_{i}^{\mathcal{D}}=d_{i}^{\mathcal{D}^{-1}}$.
The environment supports the thermodynamic cost for the change in the microscopic Boltzmann weight. Compared to eq. 3, we have modified eq. 4 by introducing a non-conservative external driving force $f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{f}_{i}^{sp}$ “along the transition,” which correspond to chemical and self-propulsion driving. The thermodynamic cost of such external driving is supported by an externally coupled reservoir of enzymatic particles (for the reactive transition) or colloidal particles (for the diffusive transitions), respectively. Since an external coupling to a reservoir is a property of a transition, $f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{f}_{i}^{sp}$ are introduced in eq. 4. Nevertheless, they could have been introduced directly in eq. 3, using exponential tilts $e^{f_{\gamma\gamma^{\prime}}^{ch}/2}$ and $e^{-f_{\gamma\gamma^{\prime}}^{ch}/2}$ for the forward and backward reactive transitions (or $e^{\vec{\mathcal{D}}\cdot\vec{f}_{i}^{sp}/2}$ and $e^{-\vec{\mathcal{D}}\cdot\vec{f}_{i}^{sp}/2}$ for the forward and backward diffusive transitions). Note that the non-reciprocal non-conservative forces $f_{i}^{nr}$ act in the state space, whereas enzymatic or self-propulsion non-conservative driving is associated with transitions between states and lies in the transition space. This fundamental physical distinction between two is clear: non-reciprocity requires the minimal notion of two states to define a dissipative current, whereas driven discrete-state processes require three states and three transitions [1].
@@
We use shorthand notations $\Delta_{\gamma\gamma^{\prime}}^{\#}(*)=(*)_{\gamma^{\prime}}^{\#}-(*)_{\gamma}^{\#}$ and $\Delta_{i}^{\vec{\mathcal{D}}\#}(*)=(*)_{i}^{\#}-(*)_{i}^{\vec{\mathcal{D}}\#}$ for the change in $\epsilon_{i}^{\#},\epsilon_{i}^{r}$, and $f_{i}^{nr}$ due to a transition. We introduce a shorthand notation $\Delta$ to refer to a transition, applied as a diffusive or reactive transition depending on the context. The set of all reactive transitions is denoted by $\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}$, and the set of all diffusive transitions is denoted by $\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}$. Four such possibilities exist for a two-dimensional square lattice: upward, downward, leftward, and rightward. Microscopic interactions and transitions are shown in fig. 2.
To demonstrate the connection of eq. 3 to the Glauber-dynamics [93] $k_{21}^{\#}=e^{\epsilon_{1}^{\#}}/(e^{\epsilon_{1}^{\#}}+e^{\epsilon_{2}^{\#}})$ and $k_{12}^{\#}=e^{\epsilon_{2}^{\#}}/(e^{\epsilon_{1}^{\#}}+e^{\epsilon_{2}^{\#}})$ or the Metropolis-dynamics [94] $k_{12}^{\#}=\min(1,e^{-\epsilon_{1}^{\#}+\epsilon_{2}^{\#}})$ and $k_{21}^{\#}=\min(1,e^{-\epsilon_{2}^{\#}+\epsilon_{1}^{\#}})$, we consider a reactive transition $\Delta_{12}^{\#}$ between type $1$ and $2$ particles with $f_{12}^{ch}=0$. Obviously, Glauber-dynamics and Metropolis-dynamics have the exponential form eq. 3, but with factors $d_{\gamma\gamma^{\prime}}=1/(e^{\epsilon_{1}^{\#}}+e^{\epsilon_{2}^{\#}})$ and $d_{\gamma\gamma^{\prime}}=1/\max{(e^{\epsilon_{1}^{\#}},e^{\epsilon_{2}^{\#}})}$, respectively, that does not represent the rate for non-interacting particles since it depends on $\epsilon_{1}^{\#}$ and $\epsilon_{2}^{\#}$. The factors non-interacting transition rates $d_{\gamma\gamma^{\prime}}$ represent the broad range of time-scales for different transitions that are present in generic reaction networks, which requires careful treatment without assumptions, see Refs.[89, 90] for details. This is carefully treated with the definition in eq. 3, which preserves the different timescales associated with different transitions. Note that the Glauber and Metropolis rates are saturating and therefore advantageous for equilibrium Monte Carlo simulations [94, 93], but they do not reflect a realistic dynamics of non-equilibrium systems. In any case the time-symmetric components of the transition rates become irrelevant after taking the ratio of transition rates for the LDB in eq. 4. The dynamics defined by eq. 3 can straightforwardly be simulated by the exact Gillespie algorithm [95].
B Multi-particle microstate transition rates.
@@
Particles with identical type and lattice index have the same dynamic and thermodynamic properties. Hence, $k_{\gamma\gamma^{\prime}}^{\#}$ are lumped together to give multi-particle transition rates $k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})=N_{\gamma^{\prime}}^{\#}k_{\gamma\gamma^{\prime}}^{\#}$ and $k_{i}^{\vec{\mathcal{D}}\#}({\{N\}})=N_{i}^{\#}k_{i}^{\vec{\mathcal{D}}\#}$. The LDB for the multi-particle transition rates (reactive ($k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})$) and diffusive ($k_{{i}}^{\vec{\mathcal{D}}\#}(\{N\})$)) is:
@@
$\begin{split}\frac{k_{\gamma\gamma^{\prime}}^{\#}({\{N\}})}{k_{\gamma^{\prime}\gamma}^{\#}({\{N+\Delta_{\gamma\gamma^{\prime}}^{\#}\}})}&=e^{-\Delta_{\gamma\gamma^{\prime}}^{\#}s^{b}+\Delta_{\gamma\gamma^{\prime}}^{\#}\epsilon+f_{\gamma\gamma^{\prime}}^{ch}},\\ \frac{k_{{i}}^{\vec{\mathcal{D}}\#}(\{N\})}{k_{i}^{(\vec{\mathcal{D}}\#)^{-1}}(\{N+\Delta_{i}^{\vec{\mathcal{D}}\#}\})}&=e^{-\Delta_{i}^{\vec{\mathcal{D}}\#}s^{b}+\Delta_{i}^{\vec{\mathcal{D}}\#}\epsilon+\vec{\mathcal{D}}\cdot\vec{f}_{i}^{sp}}.\end{split}$ (5)
@@
Here, $\Delta s^{b}$ incorporates the change in the microscopic Boltzmann entropy, defined as a global lattice property [96]:
@@
It quantifies the statistical degeneracy of the microstate and is a system property defined for the lattice. Importantly, $\Delta_{\gamma\gamma^{\prime}}^{\#}\epsilon^{r}=\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}$ and $\Delta_{i}^{\vec{\mathcal{D}}\#}\epsilon^{r}=\Delta_{i}^{\vec{\mathcal{D}}\#}\varepsilon^{int}$, as satisfied by the reciprocal microscopic Boltzmann weight. We introduce the notation $\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}=\varepsilon^{int}(\{N\})-\varepsilon^{int}(\{N+\Delta_{\gamma\gamma^{\prime}}\})$, which represents a change of a function of the lattice state, a global property; this applies similarly to $s^{b}$. In contrast, the change in the non-reciprocal microscopic Boltzmann weight $\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$ cannot be represented as a change of a global function of the lattice state. This results in a fundamental difference between the boundary ($\Delta_{\gamma\gamma^{\prime}}^{\#}\varepsilon^{int}$ and $\Delta_{\gamma\gamma^{\prime}}^{\#}s^{b}$) and the bulk terms ($\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$) of the transition affinity.